OpenStudy (mtalhahassan2):

The two equal sides of an isosceles triangle each have a length of 2x+3y-1. The perimeter of the triangle is 7x+9y. Determine the length of the third side.

OpenStudy (anonymous):

Perimeter of a triangle is the sum of the three sides. Let a represent the length of the third side. (2x + 3y - 1) + (2x + 3y -1) + a = 7x + 9y Set up the equation. 4x + 6y - 2 + a = 7x + 9y Simplify 6y - 2 + a = 3x + 9y Subtract 4x from each side. a - 2 = 3x + 3y Subtract 6y from each side. a = 3x + 3y + 2 Add 2 to each side. The length of the third side is 3x + 3y + 2.

OpenStudy (mtalhahassan2):

Thanks you so much.